OpenStudy (anonymous):

solve (5X+1)=3 round to the nearest 100th..free metal

OpenStudy (anonymous):

x=2/5

OpenStudy (anonymous):

x=.04

OpenStudy (anonymous):

i meant nearest thousandth

OpenStudy (anonymous):

(5X+1)=3 x = 3-1 x =2/5 x= 0.4 round= 0.004

OpenStudy (anonymous):

If I deserve medal then give me one